# Quantum Computing in Wireless Communications and Networking: A Tutorial-cum-Survey

### v1

@@ Abstract
Owing to its outstanding parallel computing capabilities, quantum computing (QC) has been a subject of continuous attention. With the gradual maturation of QC platforms, it has increasingly played a significant role in various fields such as transportation, pharmaceuticals, and industrial manufacturing, achieving unprecedented milestones. In modern society, wireless communication stands as an indispensable infrastructure, with its essence lying in optimization. Although artificial intelligence (AI) algorithms such as Reinforcement Learning (RL) and mathematical optimization have greatly enhanced the performance of wireless communication, the rapid attainment of optimal solutions for wireless communication problems remains an unresolved challenge. QC, however, presents a new alternative. This paper aims to elucidate the fundamentals of QC and explore its applications in wireless communications and networking. First, we will provide a tutorial on QC, covering its basics, characteristics, and popular QC algorithms. Next, we will introduce the applications of QC in communication and networking, followed by its applications in miscellaneous areas such as security and privacy, localization and tracking, and video streaming. Finally, we will discuss remaining open issues before concluding.
@@ I-A Background and Motivation
Quantum computing (QC) [1, 2], a revolutionary approach to processing information based on quantum mechanics principles, shows great potential for transforming various industries. Unlike traditional computers that use bits in a binary state (0 or 1), quantum computers apply qubits that can exist in multiple states simultaneously, allowing for much more powerful computation. This technology is expected to greatly improve cryptography by breaking current encryption methods and enabling the development of more secure ones. Additionally, QC is highly effective in solving optimization problems, which can benefit applications like route planning and managing supply chains. It also has applications in fields such as drug discovery, materials science, and simulating quantum systems, offering faster insights into complex interactions at the molecular level. As of December 2023, the investment for quantum market reaches to \$6.7B with 261 start-ups, and potential quantum technology market size by 2040 is estimated by \$173B [3], fueled by substantial investments from governments and industries harnessing the transformative potential of this cutting-edge technology.
@@
On the other hand, wireless communications and networking play a foundational role in modern connectivity, facilitating the smooth exchange of data across devices and supporting a wide range of applications, from mobile networks and WiFi to IoT devices and satellite communications [4]. Optimization is crucial in wireless networks, especially as we move towards the era of 6G [5]. With the increasing demand for higher data rates, lower latency, and greater device density, optimizing spectrum usage, reducing interference, and improving signal processing algorithms become even more critical. 6G, the next frontier in wireless technology, is poised to offer unprecedented speeds, ultra-low latency, and innovative applications [6]. Effectively optimizing these advanced wireless networks will be essential for fully realizing the potential of 6G, ensuring efficient and reliable communication in an increasingly interconnected world.
@@
Additionally, the interaction among the aforementioned enabling technologies holds the potential for significant performance enhancements, encompassing increased capacity, enhanced energy efficiency, and improved reliability. For instance, incorporating edge intelligence for micro-service orchestration, as demonstrated in [7], can bolster network throughput and reduce service delay in next-generation networks. Leveraging QC algorithms for optimization tasks can accelerate this process. As wireless networks expand to encompass intelligent devices, such as autonomous vehicles, UAVs and satellites, the complexity of problems increase accordingly [8]. In particular, optimization tasks in wireless communications and networking, like channel assignment, power allocation, edge computing, and video streaming demand substantial computational resources. While conventional solutions often rely on local servers, cloud infrastructure, or GPU servers, they still struggle to meet the growing research demands. QC techniques offer a promising approach to addressing these challenges by leveraging its ability to handle complex optimization scenarios efficiently, such as slicing [9] and Ultra-Reliable Low Latency Communications (URLLC) [10] in 6G.
@@
Given this landscape, our survey aims to explore the intersection of QC and wireless communications and networking, providing insights into the latest developments, challenges, and opportunities in this rapidly evolving field. We explore the application of QC techniques to address optimization challenges in wireless networks, offering a comprehensive overview of current research trends and future directions. Through this survey, we seek to shed light on the transformative potential of QC in reshaping the landscape of wireless communications and networking.
@@ I-B Related Surveys
While previous surveys, primarily concentrated on quantum communications [11, 12, 13], have explored the applications and advantages of integrating QC into wireless networks from various applications [14, 15, 16, 17], a comprehensive survey on QC-enabled communication and networking is still lacking. The first survey [14] investigated the utilization of QC in addressing challenges within wireless communication systems. It provided an overview of QC fundamentals using linear algebra, showing its potential for enhancing wireless communication systems as proposed in existing literature. The survey extensively explored various optimization problems encountered in both the physical and network layers of wireless communications, offering comparisons between classical and quantum-assisted solutions. However, the survey from a few years back lacked coverage of AI-powered QC algorithms such as Quantum Neural Networks (QNNs) and Quantum Deep Reinforcement Learning (QDRL). Moreover, it did not address particular concerns within wireless networks like edge computing and video streaming, which are especially relevant in the context of 5G/6G networks.
@@
AI-driven QC algorithms have been explored in several studies, including [15, 16, 17, 18]. However, in [15], the discussion on conventional and AI-driven QC algorithms remained largely conceptual, with limited elaboration on specific algorithms like QDRL. Topics such as task offloading and content caching in wireless networks were also presented conceptually. Another brief survey on QC algorithms for 6G networks, focusing on quantum-inspired machine learning applications, can be found in [16]. This survey provided a brief overview of both conventional and AI-driven algorithms, including the Shor algorithm, Grover algorithm, QRL, and QNNs. It briefly discussed a few topics mainly related to security and resource allocation for 6G networks. The fundamental principles of QC, as well as conventional and AI-driven QC algorithms, were relatively described in [17]. However, the paper provided only a limited introduction to their applications in wireless networks, focusing mainly on power allocation and channel assignment. In [18], the applications of QC were reviewed in details, with a focus on quantum cloud computing. The paper discussed the latest advancements in quantum cloud technology, covering various cloud-based models, platforms, and recently developed technologies and software use cases. It also examined several applications of quantum cloud computing, such as resource management, quantum serverless computing, and addressing security and privacy concerns. Table I presents a comparative summary of the referenced surveys, underscoring that the most recent ones fail to provide a comprehensive overview of the QC applications in wireless networks. This observation highlights a gap in the existing literature.
@@ I-C Contributions
  • We examine recent advancements in the application of QC algorithms to address various communication and networking challenges in wireless networks, spanning power allocation, channel assignment, user association, routing, edge computing/task offloading, content caching, security and privacy, localization and tracking, video streaming, and more. Additionally, we identify key challenges and open problems for future research to fully utilize the potential of QC.
@@ I-D Literature Categorization and Paper Organization
Figure 1: Paper organization
We have adopted the following methodology to ensure a comprehensive and in-depth review of applications of QC algorithms. We primarily organize the papers from two dimensions: the research issues and category of QC algorithms. We divide the research issues in wireless networks into two parts: communications from the perspective of the physical layer, where issues typically arise within a single hop, and networking from the perspective of the routing layer, where issues typically arise within multiple hops. The former mainly considers topics including channel assignment, power allocation, and user association, while the latter mainly considers routing, task offloading, content caching, and among others. For each issue, we consider its network characteristics such as the presence of multiple cells in communications and mobility in networking. Additionally, we divide QC algorithms into two categories: conventional QC algorithms, such as classical QAOA and quantum-inspired heuristics like QPSO and Quantum Genetic Algorithm (QGA), and AI-driven algorithms, such as QNNs and QDRL. The paper organization is described in Fig. 1.
@@
BS Base Station
NISQ Noisy Intermediate-Scale Quantum
@@
QPSO Quantum Particle Swarm Optimization
QML Quantum Machine Learning
QUBO Quadratic Unconstrained Binary Optimization
@@
QRL Quantum Reinforcement Learning
QDRL Quantum Deep Reinforcement Learning
QGA Quantum Genetic Algorithm
QNN Quantum Neural Network
SVM Support Vector Machine
PQC parameterized Quantum Circuits
QGLM Quantum Generative Learning Model
NOMA Non-orthogonal Channel Multiple Access
MIMO Multiple Input Multiple Output Antennas
FDMA Frequency-division Multiple Access
OFDMA Orthogonal Frequency-division Multiple Access
RSU Road-side Unit
IRS Intelligent Reflecting Surface
@@
VQE Variational Quantum Eigensolver
QMARL Quantum Multiple Agent Reinforcement Learning
ILP Integer Linear Programming
QEA Quantum Evolution Algorithm
URLLC Ultra-Reliable Low Latency Communications
@@ II-A The Postulates of Quantum Mechanics
QC is grounded in the principles of quantum mechanics and unequivocally adheres to its fundamental rules. In this section, we will outline the essential principles of quantum mechanics [19, 20], providing the requisite foundation for the subsequent discussion. This explanation is crucial in establishing a mathematical framework for comprehending QC, thereby facilitating a deeper understanding of this intricate field for readers without a background in physics.
@@
The Hilbert space, as delineated in the given postulate, corresponds to a complete complex linear space, characterized by a defined inner product, pertinent to QC since it does not involve infinite spatial dimensions. A collection of $n$ linearly independent vectors is capable of spanning an n-dimensional linear space. When these vectors are both normalized and mutually orthogonal, they constitute what is known as orthonormal bases for the aforementioned space. A unique unit vector, designated as a ’state’, encapsulates the description of a specific physical system. To articulate this with greater precision and mathematical rigor, the state of a physical system is depicted by a unit vector. This unit vector can be linearly decomposed using any complete set of orthonormal bases within the linear space, offering a mathematically precise representation of the physical system in question.
Physicist Dirac creates a set of symbols to describe system states [21]. The symbol $\ket{.}$ is called ket, corresponding to a column unit vector in the space. Thus, a specific system state, denoted by a quantum state |$\ket{\phi}$ , can be represented linearly by a set of orthonormal bases {|$\{\ket{a_{1}},\ket{a_{2}},\dots,\ket{a_{n}}\}$.
|$\ket{\phi}=\sum_{i=1}^{n}c_{i}\ket{a_{i}},\quad st.\sum_{i}|c_{i}|^{2}=1.$ (1)
The algebra reveals that the orthonormal bases of a Hilbert space are not unique; thereby, a quantum state can be described by different bases. This property is known as representation transformation in quantum mechanics and is essentially equivalent to the basis transformation.
@@
In general, finding the Hamiltonian to characterize a particular physical system is difficult (requiring substantial experimental results), but fortunately this is not the primary concern for the involved theoretical framework. In the discussions of QC and quantum information, we always assume that $H$ is known. Thus, if the system in the state of time $t_{0}$ is |$\ket{\phi(t_{0})}$, then the system in the $t_{1}$ moment can be solved by the Schrodinger equation in eq. (3),
@@
where $U(t)$ is the evolutive operator from $t_{0}$ to $t_{1}$, defined as:
@@
Assuming that the state of the system before measurement is |$\ket{\phi}$, the probability of observing the result $m$ is in eq. (6),
@@
The post-measurement state is in eq. (7),
@@
If the index number of subsystems is $1$ to $n$, and the state of subsystem-$i$ can be represented as |$\ket{\phi_{i}}$, then the state of the entire system is |$\ket{\phi_{1}}\otimes\cdots\otimes\ket{\phi_{n}}$. The notation $\otimes$ represents the tensor product, which is a way of combining smaller vector spaces together to form a larger vector space. Its operation rules can be denoted by Kronecker multiplication in eq. (8),
@@
where A and B are $m\times n$ and $p\times q$ matrix, respectively.
@@ II-B The Characteristics of QC
In the domain of classical computing, the bit, with possible states 00 or $1$, serves as the fundamental unit of information. This binary state system can represent a physical system with two clearly distinguishable states, exemplified by the widely utilized high and low-level electrical signals.
Conversely, in QC, the basic information unit is the quantum bit (qubit). A qubit has the capability to physically manifest in a variety of forms. Examples of this include the ground state and the first excited state of an electron within a hydrogen atom; the $+\frac{1}{2}$ and $-\frac{1}{2}$ components of a proton’s spin in any given direction; or the clockwise and anti-clockwise circular polarization of light. A notable distinction of qubits, as opposed to classical bits, is their ability to exist not merely in a binary state of ‘dimorphism’ but also in any intermediary phase between these two states, a phenomenon known as linear superposition. Adhering to Postulate 1, a qubit can be mathematically represented as a unit vector within a two-dimensional Hilbert space, as delineated in eq. (9)
|$\ket{\phi}=\alpha\ket{0}+\beta\ket{1}\quad st.\ |\alpha|^{2}+|\beta|^{2}=1,$ (9)
where the complex numbers $\alpha$ and $\beta$ satisfy the normalization condition, and |$\ket{0}$ and |$\ket{1}$ are orthonormal basis $\begin{pmatrix}1\\ 0\end{pmatrix}$ and $\begin{pmatrix}0\\ 1\end{pmatrix}$. Thereby, |$\ket{\phi}$ is represented as $\begin{pmatrix}\alpha\\ \beta\end{pmatrix}$.
@@
A qubit is a unit vector in a two-dimensional Hilbert space. The representation of two qubits can be derived by (Kronecker product eq. 8) of two single qubits. For example, |$\ket{\phi_{1}}=\alpha_{1}\ket{0}+\beta_{1}\ket{1}$, |$\ket{\phi_{2}}=\alpha_{2}\ket{0}+\beta_{2}\ket{1}$, then two qubits:
@@
where |$\ket{00}=\begin{pmatrix}1\\ 0\\ 0\\ 0\end{pmatrix}$, |$\ket{01}=\begin{pmatrix}0\\ 1\\ 0\\ 0\end{pmatrix}$, |$\ket{10}=\begin{pmatrix}0\\ 0\\ 1\\ 0\end{pmatrix}$, and |$\ket{11}=\begin{pmatrix}0\\ 0\\ 0\\ 1\end{pmatrix}$, form a set of orthonormal bases for a four-dimensional Hilbert space. Obviously, $|\alpha_{1}\alpha_{2}|^{2}+|\alpha_{1}\beta_{2}|^{2}+|\beta_{1}\alpha_{2}|^{2}+|\beta_{1}\beta_{2}|^{2}=1$, and |$\ket{\phi_{1}\phi_{2}}$ is a normalized vector in a four-dimensional Hilbert space. By analogy, the extension of $n$ qubits |$\ket{\phi_{1}\phi_{2}\cdots\phi_{n}}$ can be expressed in eq. (11),
|$\ket{\phi_{1}\phi_{2}\cdots\phi_{n}}=\sum_{i=0}^{2^{n}-1}c_{i}\ket{i}\qquad st.\ \sum|c_{i}|^{2}=1,$ (11)
Suppose a quantum register in a quantum computer stores $n$ qubits. According to eq. (10), the quantum register stores every number from 00 to $12^{n}-$ 1 simultaneously with the probability of $|c_{i}|^{2}$. In other words, a $n$-qubit quantum register holds $2^{n}$ binary numbers.
@@
While $n$ superposed qubits can concurrently represent $2^{n}$ binary numbers, there is generally no methodology to simultaneously extract each binary number concealed within the superposition state. More specifically, one cannot accurately determine the probability amplitude $c_{i}$ (eq. 11) of each individual binary number for an unknown $n$-qubit. Our primary tool for understanding an unknown quantum state is through measurement or observation. However, each observation only yields a fragment of information about the initial state since the state collapses upon measurement, leading to significant data loss.
To illustrate, consider a beam of light emitting a mixture of horizontally and vertically polarized photons |$\ket{\phi}=\alpha\ket{\uparrow}+\beta\ket{\rightarrow}$. Once this beam passes through a horizontal polarization filter, only vertically polarized photons $\ket{\uparrow}$ can be observed. This observation results in |ϕ|ketitalic-$\ket{\phi}\Rightarrow\ket{\uparrow}$, with the frequency of observing vertically polarized light (or vertically polarized photons) approximating $|\alpha|^{2}$. Similarly, if this light is subjected to a vertical polarization filter, the observable light will be horizontally polarized with a probability of $|\beta|^{2}$.
The methodology of this aforementioned observation characterizes the process of projective measurement, a specific case of the more general measurement process described by Postulate 3. This type of measurement, alternatively known as Von Neumann measurement, is extensively employed across various studies. Regardless of physical circumstances, this can be abstracted as an observable or a Hermitian operator $M$, presented in the spectral decomposition form as shown below.
@@
where the projective operator $P_{m}$ is the projection onto the eigenspace $M_{m}$ corresponding to the eigenvalue $m$, $P_{m}P_{m^{\prime}}=\delta_{mm^{\prime}}P_{m}$, and satisfies completeness condition $\sum_{m}P_{m}=I$.
The measurement outputs are eigenvalues $m$. When measuring the state |$\ket{\phi}$, the probability of getting a certain $m$ is in eq. (13),
@@
After the measurement, |$\ket{\phi}$ collapses to eq. (14),

### v3

@@ Abstract
Owing to its outstanding parallel computing capabilities, quantum computing (QC) has been a subject of continuous attention. With the gradual maturation of QC platforms, it has increasingly played a significant role in various fields such as transportation, pharmaceuticals, and industrial manufacturing, achieving unprecedented milestones. In modern society, wireless communication stands as an indispensable infrastructure, with its essence lying in optimization. Although artificial intelligence (AI) algorithms such as reinforcement learning (RL) and mathematical optimization have greatly enhanced the performance of wireless communication, the rapid attainment of optimal solutions for wireless communication problems remains an unresolved challenge. QC, however, presents a new alternative. This paper aims to elucidate the fundamentals of QC and explore its applications in wireless communications and networking. First, we provide a tutorial on QC, covering its basics, characteristics, and popular QC algorithms. Next, we introduce the applications of QC in communications and networking, followed by its applications in miscellaneous areas such as security and privacy, localization and tracking, and video streaming. Finally, we discuss remaining open issues before concluding.
@@ I-A Background and Motivation
QC [1, 2], a revolutionary approach to processing information based on quantum mechanics principles, shows great potential for transforming various industries. Unlike traditional computers that use bits in a binary state (0 or 1), quantum computers apply qubits that can exist in multiple states simultaneously, allowing for much more powerful computation. This technology is expected to greatly improve cryptography by breaking current encryption methods and enabling the development of more secure ones. Additionally, QC is highly effective in solving optimization problems, which can benefit applications like route planning and managing supply chains. It also has applications in fields such as drug discovery, materials science, and simulating quantum systems, offering faster insights into complex interactions at the molecular level. As of December 2023, the global quantum market has reached an investment total of \$ 6.7 billion, with 261 active start-ups. The potential market size for quantum technology by 2040 is projected to reach \$ 173 billion [3], driven by significant investments from governments and industries aiming to leverage the transformative capabilities of quantum technologies.
@@
On the other hand, wireless communications and networking play a foundational role in modern connectivity, facilitating the smooth exchange of data across devices and supporting a wide range of applications, from mobile networks and WiFi to IoT devices and satellite communications [4]. Optimization is crucial in wireless networks, especially as we move towards the era of 6G [5]. With the increasing demand for higher data rates, lower latency, and greater device density, optimizing spectrum usage, reducing interference, and improving signal processing algorithms become even more critical. 6G, the next frontier in wireless technology, is expected to offer unprecedented speeds, ultra-low latency, and innovative applications [6]. Effectively optimizing these advanced wireless networks will be essential for fully realizing the potential of 6G, ensuring efficient and reliable communication in an increasingly interconnected world.
@@
Additionally, the interaction among the aforementioned enabling technologies holds the potential for significant performance enhancements, encompassing increased capacity, enhanced energy efficiency, and improved reliability. For instance, incorporating edge intelligence for micro-service orchestration, as demonstrated in [7], can bolster network throughput and reduce service delay in next-generation networks. Leveraging QC algorithms for optimization tasks can accelerate this process. As wireless networks expand to encompass intelligent devices, such as autonomous vehicles, UAVs and satellites, the complexity of problems increases accordingly [8]. In particular, optimization tasks in wireless communications and networking, like channel assignment, power allocation, edge computing, and video streaming demand substantial computational resources. While conventional solutions often rely on local servers, cloud infrastructure, or GPU servers, they still struggle to meet the growing research demands. QC techniques offer a promising approach to addressing these challenges by leveraging its ability to handle complex optimization scenarios efficiently, such as slicing [9] and ultra-reliable low latency communications (URLLC) [10] in 6G.
@@
Given this landscape, our survey aims to explore the intersection of QC and wireless communications and networking, providing insights into the latest developments, challenges, and opportunities in this rapidly evolving field. We explore the application of QC techniques to address optimization challenges in wireless networks, offering a comprehensive overview of current research trends and future directions. Note that while quantum key distribution is indeed a crucial area in quantum communications, this survey focuses specifically on QC applications in wireless communications and networking.
@@ I-B Related Surveys
While previous surveys, primarily concentrated on quantum communications [11, 12, 13], have explored the applications and advantages of integrating QC into wireless networks from various applications [14, 15, 16, 17], a comprehensive survey on QC-enabled communications and networking is still lacking. The first survey [14] investigates the utilization of QC in addressing challenges within wireless communication systems. It provides an overview of QC fundamentals using linear algebra, showing its potential for enhancing wireless communication systems in existing literature. The survey extensively explores various optimization problems encountered in both the physical and network layers of wireless communications, offering comparisons between classical and quantum-assisted solutions. However, the survey from a few years back lacks coverage of AI-powered QC algorithms such as quantum neural networks (QNNs) and quantum deep reinforcement learning (QDRL). Moreover, it does not address particular concerns within wireless networks like edge computing and video streaming that are especially relevant in the context of 5G/6G networks.
@@
AI-driven QC algorithms have been explored in several studies, including [15, 16, 17, 18]. However, in [15], the discussion on conventional and AI-driven QC algorithms remains largely conceptual, with limited elaboration on specific algorithms like QDRL. Topics such as task offloading and content caching in wireless networks are also presented conceptually. Another brief survey on QC algorithms for 6G networks, focusing on quantum-inspired machine learning (ML) applications, can be found in [16]. This survey provides a brief overview of both conventional and AI-driven algorithms, including the Shor algorithm, Grover algorithm, quantum reinforcement learning (QRL), and QNNs. It briefly discusses a few topics mainly related to security and resource allocation for 6G networks. The fundamental principles of QC, as well as conventional and AI-driven QC algorithms, are relatively described in [17]. However, the paper provides only a limited introduction to their applications in wireless networks, focusing mainly on power allocation and channel assignment. In [18], the applications of QC are reviewed in detail, with a focus on quantum cloud computing. The paper discusses the latest advancements in quantum cloud technology, covering various cloud-based models, platforms, and recently developed technologies and software user cases. It also examines several applications of quantum cloud computing, such as resource management, quantum serverless computing, and addressing security and privacy concerns. Table I presents a comparative summary of the referenced surveys, underscoring that the most recent ones fail to provide a comprehensive overview of the QC applications in wireless networks. This observation highlights a gap in the existing literature.
@@ I-C Contributions
  • We examine recent advancements in the application of QC algorithms to address various communication and networking challenges in wireless networks, spanning power allocation, channel assignment, user association, routing, edge computing/task offloading, content caching, security and privacy, localization and tracking, video streaming, and so forth. Additionally, we identify key challenges and open problems for future research to fully utilize the potential of QC.
@@ I-D Literature Categorization and Paper Organization
Figure 1: Paper organization.
We have adopted the following methodology to ensure a comprehensive and in-depth review of applications of QC algorithms. We primarily organize the papers from two dimensions: the research issues and category of QC algorithms. We divide the research issues in wireless networks into two parts: communications from the perspective of the physical layer, where issues typically arise within a single hop, and networking from the perspective of the routing layer, where issues typically arise within multiple hops. The former mainly considers topics including channel assignment, power allocation, and user association, while the latter mainly considers routing, task offloading, content caching, and among others. For each issue, we consider its network characteristics such as the presence of multiple cells in communications and mobility in networking. Additionally, we divide QC algorithms into two categories. One is conventional QC algorithms, such as classical QAOA, quantum-inspired heuristics like QPSO, and quantum genetic algorithm (QGA). The other is AI-driven algorithms, such as QNNs and QDRL. The paper organization is described in Fig. 1.
@@
URLLC ultra-reliable low latency communications
QNN quantum neural network
QDRL quantum deep reinforcement learning
QRL quantum reinforcement learning
@@
QPSO quantum particle swarm optimization
QGA quantum genetic algorithm
NISQ noisy intermediate-scale quantum
VQE variational quantum eigensolver
QML quantum machine learning
PQC parameterized quantum circuits
QUBO quadratic unconstrained binary optimization
QGLM quantum generative learning model
BS base station
NOMA non-orthogonal channel multiple access
OFDMA orthogonal frequency-division multiple access
MIMO multiple input multiple output antennas
@@
QMARL quantum multiple agent reinforcement learning
RSU road-side unit
@@
IRS intelligent reflective surface
QEA quantum evolution algorithm
ILP integer linear programming
@@ II A Tutorial of QC
This section is organized into three parts. The first part introduces the fundamental principles and characteristics of quantum algorithms (Subsections A-D). The second part examines prominent quantum algorithms that are expected to achieve quantum supremacy during the noisy intermediate-scale quantum (NISQ) era (Subsections E and F). The final part covers widely used quantum hardware and software platforms. These sections are structured sequentially to support the discussion of quantum computing (QC) applications in wireless networks.
@@ II-A The Postulates of Quantum Mechanics
QC is grounded in the principles of quantum mechanics. In this section, we outline the essential principles of quantum mechanics, providing the requisite foundation for the subsequent discussions. This explanation is crucial in establishing a mathematical framework for comprehending QC, thereby facilitating a deeper understanding of this field for readers without a background in physics.
@@
The Hilbert space, as described in Postulate 1, is a complete complex linear space characterized by a defined inner product, relevant to QC since it does not involve infinite spatial dimensions. A collection of $n$ linearly independent vectors is capable of spanning an $n$-dimensional linear space. When these vectors are both normalized and mutually orthogonal, they form orthonormal bases for the given space. A unique unit vector, referred to as a ‘state,’ encapsulates the description of a specific physical system. This unit vector can be linearly decomposed using any complete set of orthonormal bases within the linear space, offering a mathematically precise representation of the physical system.
Physicist Dirac creates a set of symbols to describe system states [19]. The symbol $\ket{\bullet}$ is termed a ket, corresponding to a column unit vector in the space. Thus, a specific system state, denoted by a quantum state |$\ket{\phi}$ in eq. (1), can be represented linearly by a set of orthonormal bases {|$\{\ket{a_{1}},\ket{a_{2}},\dots,\ket{a_{n}}\}$.
|$\ket{\phi}=\sum_{i=1}^{n}c_{i}\ket{a_{i}},\quad\text{s.t.}\sum_{i}|c_{i}|^{2}=1.$ (1)
The algebra reveals that the orthonormal bases of a Hilbert space are not unique; thereby, a quantum state can be described by different bases. This property is known as representation transformation in quantum mechanics, which is essentially equivalent to basis transformation.
@@
In general, finding the Hamiltonian to characterize a particular physical system is difficult (requiring substantial experimental results), but fortunately this is not the primary concern for the involved theoretical framework. In the discussions of QC and quantum information, we always assume that $H$ is known. Thus, if the system in the state at time $t_{0}$ is |$\ket{\phi(t_{0})}$, then the system at $t_{1}$ can be solved by the Schrodinger equation in eq. (3),
@@
where $U(t)$ is an evolutive operator from $t_{0}$ to $t_{1}$ in eq. (4),
@@
Assuming that the state of the system before measurement is |$\ket{\phi}$, the probability of observing the result $m$ is given by eq. (6),
@@
The post-measurement state is given by eq. (7),
@@
If the index numbers of the subsystems range from $1$ to $n$, and the state of subsystem-$i$ can be represented as |$\ket{\phi_{i}}$, then the state of the entire system is |$\ket{\phi_{1}}\otimes\cdots\otimes\ket{\phi_{n}}$. The notation $\otimes$ represents a tensor product, which combines smaller vector spaces to form a larger vector space. Eq. (8) presents general rules for the tensor product, also known as the Kronecker product,
@@
where A and B are $m\times n$ and $p\times q$ matrix, respectively. Since a quantum state is essentially an $n\times 1$ matrix, eq. (8) is applicable. If quantum states are described by density operators (see ”D. The Description by Density Operators”), eq. (8) can be directly applied.
@@ II-B The Characteristics of QC
The main advantage of QC arises from quantum parallelism, which is rooted in the unique properties of qubits.
@@
In classical computing, the bit, which can have a value of either 00 or $1$, serves as the fundamental unit of information. This binary state system can represent a physical system with two clearly distinguishable states, such as the widely used high and low-level electrical signals.
The fundamental unit of information in QC is the quantum bit, or qubit. Unlike classical bits, which exist strictly in one of two states (00 or $1$), qubits can assume multiple physical forms. Examples include the ground and excited states of an electron in a hydrogen atom, the spin states of a proton ($+\frac{1}{2}$ and $-\frac{1}{2}$), or the clockwise and counterclockwise circular polarization of light. A key distinction of qubits is their ability to exist in a superposition, where they can simultaneously represent both 00 and $1$, as well as any linear combination of these states. Mathematically, a qubit is represented as a unit vector in a two-dimensional Hilbert space, as described by Postulate 1 and shown in eq. (9).
|$\ket{\phi}=\alpha\ket{0}+\beta\ket{1},\quad\text{s.t.}|\alpha|^{2}+|\beta|^{2}=1,$ (9)
[…]
@@
A qubit is a unit vector in a two-dimensional Hilbert space. The representation of two qubits can be derived by (Kronecker product eq. 8) two single qubits. For example, |$\ket{\phi_{1}}=\alpha_{1}\ket{0}+\beta_{1}\ket{1}$, |$\ket{\phi_{2}}=\alpha_{2}\ket{0}+\beta_{2}\ket{1}$, then two qubits:
@@
[…]
|$\ket{\phi_{1}\phi_{2}\cdots\phi_{n}}=\sum_{i=0}^{2^{n}-1}c_{i}\ket{i},\qquad\text{s.t.}\sum|c_{i}|^{2}=1.$ (11)
Suppose a quantum register in a quantum computer stores $n$ qubits. According to eq. (10), the quantum register stores every number from 00 to $12^{n}-$ 1 simultaneously with the probability of $|c_{i}|^{2}$. In other words, an $n$-qubit quantum register holds $2^{n}$ binary numbers.
@@
While $n$ superposed qubits can concurrently represent $2^{n}$ binary numbers, there is generally no method to extract each binary number concealed within the superposition state simultaneously. More specifically, it is impossible to accurately determine the probability amplitude $c_{i}$ in eq. 11 of each individual binary number for an unknown $n$-qubit. The primary tool for probing an unknown quantum state is through measurement or observation. However, each observation only yields a fragment of information about the original state since the state collapses upon measurement, leading to significant data loss.
To illustrate, consider a beam of light emitting a mixture of horizontally and vertically polarized photons |$\ket{\phi}=\alpha\ket{\uparrow}+\beta\ket{\rightarrow}$. Once this beam passes through a horizontal polarization filter, only vertically polarized photons $\ket{\uparrow}$ can be observed, resulting in |ϕ|ketitalic-$\ket{\phi}\Rightarrow\ket{\uparrow}$, with the frequency of observing vertically polarized light (or vertically polarized photons) approximating $|\alpha|^{2}$. Conversely, if this light is subjected to a vertical polarization filter, the observable light is horizontally polarized with a probability of $|\beta|^{2}$.
The methodology of this aforementioned observation characterizes the process of projective measurement, a specific case of the more general measurement process described by Postulate 3. This type of measurement, also known as Von Neumann measurement, is extensively employed across various studies. Regardless of physical circumstances, this can be abstracted as an observable or a Hermitian operator $M$, presented in the spectral decomposition form in eq. 12,
@@
where the projective operator $P_{m}$ is the projection onto the subspace of $M$ corresponding to the eigenvalue $m$. It satisfies $P_{m}P_{m^{\prime}}=\delta_{mm^{\prime}}P_{m}$ and the completeness condition $\sum_{m}P_{m}=I$.
The measurement outputs are eigenvalues $m$. When measuring the state |$\ket{\phi}$, the probability of getting a certain $m$ is given by:
@@
After the measurement, |$\ket{\phi}$ collapses to: